OpenStudy (anonymous):

How to put y=4cos(x) in X= form?

OpenStudy (anonymous):

\[x=\cos^{-1} (y/4)\]

OpenStudy (anonymous):

It is only allowed in said domain in order to make it the inverse function (It needs to pass the horizontal line test)

OpenStudy (anonymous):

Inverse y = 4cos(x) x/4 = 4cos(x)/4 y/4 = cos(x) take inverse of each side arccos(y/4) = arccos(cos(x)) arccos(y/4) = 1x arccos(y/4) = x Remember that arccos is the inverse function of cos(x) and is limited to the domain [0, pi]

OpenStudy (anonymous):

ignore the mistake i made in the variables

OpenStudy (anonymous):

Here is the problem I'm working at:

OpenStudy (anonymous):

Find the volume of the solid generated by revolving the described region about the given axis: The region in the first quadrant bounded above by the line y=4 and by the curve y=4sin(x) for the interval 0≤x≤π2 about the line y=4

OpenStudy (anonymous):

I think I'm using the cross-section method, but am not too sure on the radius.

OpenStudy (anonymous):

Which I think would be (4-4sin(x)

OpenStudy (anonymous):

Wait do you want 4sin(x) to equal 4?

OpenStudy (anonymous):

I'm honestly not too sure on where to start with this problem.

OpenStudy (anonymous):

I'm confused by your question but if you want it to equal four you can use the unit circle and think where is cos(x) = 1, the answer being pi

OpenStudy (anonymous):

Are you familiar with solids of rotation?

OpenStudy (anonymous):

tbh no I know trig functions though but meh you should ask in chat for help

OpenStudy (anonymous):

haha alright, thanks for the help

OpenStudy (anonymous):

@Cinar, do you know which method I would use here?

OpenStudy (anonymous):

little bit (: I am trying to find it

OpenStudy (anonymous):

Sweet, thanks

OpenStudy (anonymous):

what is the rotation axis?

OpenStudy (anonymous):

it is y=4

OpenStudy (anonymous):

so I'm thinking the radius is (4-4sin(x))

OpenStudy (anonymous):

Any luck?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[V=\pi \int\limits_{0}^{2\pi}(4-4\sin x)^2dx\]

OpenStudy (anonymous):